OpenStudy (anonymous):

Which of the following represents fifth root of x squared in exponential form?

OpenStudy (anonymous):

@Nnesha

OpenStudy (zzr0ck3r):

\(\sqrt[a]{x^b}=x^{\frac{b}{a}}\)